1. Eddie borrows R12000 from Janet at the beginning of 2005, at a rate of 10% p.a. compound interest. At the end of each year, he pays back R500. How much must he still pay at the end of 2008? How much does he have to pay back altogether?
1
Expert's answer
2022-03-10T15:18:40-0500

annual payment

"A=12000(\frac{0.1(1+0.1)^4}{(1+0.1)^4-1})=3785.65"

"3785.65\times4-500\times4=13142.60"

"3785.65\times4=15142.60" much does he have to pay back